OpenStudy (anonymous):

So I have a probability function that satisfies the following: P(0)=P(1) and P(k+1)=(1/k)P(k) for k = 1, 2, 3... Okay... so I know that P(0)=P(1)=P(2) as well as P(3)=(1/2)P(2), P(4)=(1/3)P(3), ..... So my question is... how do I find P(k)?

OpenStudy (anonymous):

The whole point of this is that I am trying to solve for P(0)... I know that \[\sum_{i=0}^{\infty}P(i)=1\] So therefore \[P(0)+\sum_{k=1}^{\infty}P(k) = 1\] So I really really really need to find that P(k)... pretty please with sugar on top someone...... anyone........ CRICKETS!!!!

OpenStudy (anonymous):

Maybe Zale is on a snack break lol.

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

Or if maybe anyone has a strategy on how to work backwards on this type of problem that might at least help a little... answer choices are A) ln e B) e - 1 C) (e+1)^-1 D) e^-1 E) (e-1)^-1

OpenStudy (zarkon):

\[p(1)=p(0)\] \[p(2)=\frac{p(1)}{1}=p(0)\] \[p(3)=\frac{p(2)}{2}=\frac{p(0)}{2!}\] \[p(4)=\frac{p(3)}{3}=\frac{p(0)}{3!}\] ... \[p(n)=\frac{p(0)}{(n-1)!}\]

OpenStudy (anonymous):

So then \[P(0)+\sum_{k=1}^{\infty}\frac{ P(0) }{ (k-1)! }=1\] \[P(0)+P(0)\sum_{k=1}^{\infty}\frac{ 1 }{ (k-1)! }=1\] \[P(0)+P(0)e=1\] \[P(0)(1+e)=1\] \[P(0)=\frac{ 1 }{ (1+e) }\]

OpenStudy (anonymous):

I understand how you got p(k) by looking at yours but even me staring at the thing for hours I would of never been able to put that together... :/ thank you for the help though. I appreciate it.